OpenStudy (anonymous):

A wire is stretched by 8 mm when a load of 60 N is applied. What will be the extension of a wire of the same material having four times the cross-sectional area and twice the original length, when the same load is applied?

OpenStudy (anonymous):

Hooke's law is \[F = kx\] If you know F = 60 N and x = 8 mm, then you can calculate the spring constant k. The spring constant is related to Young's modulus E by \[k = \frac{AE}L\] where A is cross sectional area and L is length. The value of E is inherent to a material and does not depend on the geometry of the wire/spring/etc. Hence \[k = \frac{F}x \propto \frac{A}L\]